North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6

Wednesday 14th March 2018


A pleasant, sunny at times, Spring like day at Grosmont yesterday where even the mud in the MPD yard was beginning to dry up. In attendance were Bill Dobson, Ian McCall, Steve Hyman, Derek Shorten, Nigel Hall, Jon Bradley and myself. Fortunately no mice as Bill had forgotten to bring some new mouse traps.

After the morning tea, Derek got on with painting the steam heat and vacuum pipes under the tender with a coat of black gloss, and Nigel made a start on prepping the J27 tender buffer beam ready for John Furness: it is now in anti corrosion green primer. I spent the morning at the MPD checking on progress with the J27, about which a separate note will follow. Ian McCall, on his last visit for some weeks, machined down the eccentric locking nuts on the lathe in the workshop. The rest of the team put up the inside left eccentric, having decided that this would not prevent the outside one being fitted once it had been straightened. Lunch was had on the bench in the Shed.

After lunch, Barney came up to Deviation Shed to check the eccentrics, and stayed to help, with advice and muscle power, the fitting of the left hand connecting rod. This involved moving the locomotive forward half a turn to get the big end down to the bottom so the rod would clear the step, and, with a bit of jiggling, and a lot of heaving, the rod went home in the crosshead. In the absence of the big end bearing, a block of wood was put in to keep the rod off the bearing axle. The bearing itself remains at the MPD awaiting machining after being remetalled. Steve then spent the rest of the afternoon securing the little big end with its nut and driving the cotter home - which also involved removing and refitting the recalcitrant oil pot to get access. In the meantime, the by now straightened, skimmed and cooled, left outside eccentric rod was returned to Deviation Shed, was satisfactorily NDT'd during the afternoon tea break (yes, we have three tea and biscuit breaks in total during the day to keep us going!) and then Jon, Bill and myself fitted it to the locomotive. All the nuts are only finger tight, and the securing pin in the plug at the valve end was not fitted. All the eccentrics will need a thorough check for nuts, pins and cotters. In the meantime, Ian was making new cotters for the eccentrics, but the pillar drill failed and a number will need the split pin holes drilling using the hand drill unless the pillar drill can be cajoled back into operation.

There were only a small number of visitors, but the B1 was still in the MPD, having been stopped on Saturday from going to the Severn Valley because of problems with its springs, discovered during efforts to weigh it. There remains some doubt as to whether it will make its GB Scottish railtour participation, although every effort is being made to get it there. In any case, the Esk Valley line is closed following a bridge strike at Castleton. Initially thought to be by a bin lorry, it now appears it could have been a Yorkshire Water vehicle. Either way, the strike was such that the bridge has apparently moved 10 inches on the stone piers and the track has been deflected as a result. It is hoped repairs will be completed by tomorrow, but it meant that Tornado, which was also due to go to the Severn Valley yesterday, had to go down to New Bridge, and then be taken by road. There are also reported to be problems with the firebox of Eric Treacy which mean it may not be returning to traffic any time soon.

Although there is still work that can be done on the Q6, we are now waiting for the MPD team to complete a number of jobs before we can finish the bottom end and get on with the boiler retubing. With a number of individuals also unavailable this weekend, it has therefore been decided to have a break, and there will be no working party this Saturday. The next one will therefore be on Wednesday 21 March. Hope you will be there.


Saturday 10th March 2018


Ian (who was present following a cancelled diesel course), reports a murky, wet, morning at Grosmont with Andy and Joan Lowes, Peter Ellis, Trevor Wilford, and Bryan Orange and the JV's. No mice to report again but he thinks new traps are required as the chocolate has been consumed from our trap - can Bill bring some new traps please! There was fantastic support from some of the JV's parents, with Ethan Humble's Mum and Grandmother manning a sales stall, with cakes, tea and coffee. Another JV's Dad, Andy Pritchard, set up a small model railway layout alongside the stalls. These all attracted lots of visitors into Deviation Shed and also watching Tornado passing by on its way to Pickering, the B1 shunting around the yard and steam heating stock at the Station, and Standard Tank 80136 on passenger work. About £110 was made at the stalls - a tremendous effort and much appreciated by the Group.

Some JVs were working on the J27, some with Piglet working on Lucie, and some working on the Q6. Both valve bores were cleaned with fine emery paper by one of our JVs, with Joan Lowes supervising other JVs on other work, including full greasing of the locomotive and tender. One new JV was working with Bryan, refitting plugs back on to an extension cable and light fitting which got damaged last week. Peter spent the day making new cotters for the eccentrics, and has made and fitted cotters to the left hand side eccentrics. Some of these will need the locking nut reducing slightly to make a better fit, and hopefully Ian McCall will be able to operate the lathe in our workshop to do this on Wednesday. Trevor removed an old stud which was jammed in the regulator valve and fitted two new studs. This valve has now been put back upstairs on the shelf. Andy, with one of the JV's in attendance, worked on each valve in the workshop, completing the assembly with cotter pins ready for fitting. Ian has emphasisied that if the valves have to be dismantled for any reason, the new keys which Andy fitted inside at the rear end of the valve have a small hole tapped into them. An M4 bolt can be screwed into the key for easy removal. Ian himself was working fitting taper pins into the front end of the eccentric rods: a long slow process getting pins from the shed, cutting them to size and filling and splitting. The top two are still to do on Wednesday.

Lunch was had down at the MPD mess room. Afterwards, Andy, Trevor and JV Henry Pritchard fitted the valves. These are still to connect to the valve rods, but the cotters seem to be too loose when fitted and it looks as if new ones will have to be made: the old ones are on the pit wall below each valve. Two strops were found by Trevor on the engine which seem to have missed the inspection a couple of weeks ago. These are 2metre I ton strops Nos 856525 and 856527 and they have been put upstairs in the metal filling cabinet, 3rd drawer down . If so they will have to be inspected next time there is an inspection about 6 months time but in the meantime do not use these strops until their status has been clarified.

Trevor will not be available for the next three Saturdays, and next Wednesday is the last that Ian McCall will be available for six weeks.

The next working Party will be on Wednesday (14th March), by when, hopefully the last of the flue tube holes will have been chamfered so work on the boiler can be progressed.

Last Updated on Thursday, 15 March 2018 11:04

Wednesday 7th March 2018


A much better day weather wise at Grosmont yesterday - a lot milder than of late and glorious sunshine in the middle of the day as Tornado passed Deviation Shed. Tornado also worked the 09.30 train to Pickering, as well as the lunchtime departure and a special train for A1 Trust covenantors and supporters at 15.40, while Repton took the 10.30 train to Pickering and 80136 was pottering up and down the line to Goathland on running in turns. What more could you wish for? (but don't answer that question!). There was a really big turn out at Deviation Shed with 10 of us, later being joined by two young cleaners from the MPD. To mark the occasion, Daniel Gibbard took the attached photo of the assembled miltitude at lunchtime, but for the record those present were Bill Dobson, Ian McCall, Derek Shorten, Steve Hyman, Nigel Hall, Roy Marshall, Jon Bradley, Dave Donegan, Ian Pearson and myself, joined by the two cleaners - Chris Wakefield and Daniel Gibbard. But no mice. Fortunately there was plenty of milk, so the kettle was on almost constantly to provide cups of tea for such a gathering before getting to work. We might need a bigger teapot if the numbers continue at this level!

Q6 volunteer working partu 7th March - Daniel Gibbard

Derek put a coat of paint on the vacuum and steam heat pipes under the tender and cab, as well as washing up and making some of the many brews. The two Ians, with assistance from Jon, removed the damaged left hand forward eccentric rod which was taken down at the MPD for inspection and rectification. There was much scratching of heads over the best way to effect the repair but they were due to give it a go today. The Ians then fitted the right hand eccentric rods, but note, these are only fastened hand tight and the taper pins at the front of the rod are just loosely fitted. Roy spent most of the day making cotters for the eccentrics. Bill, Dave, Steve and Nigel fitted the crosshead to the piston rod on both sides of the locomotive.

Three roe deer grase the bank above Armstrong Oilers - Nigel Hall

Lunch was taken in Deviation shed while our two young cleaners volunteered to go into the boiler to fit the regulator rod through the retaining bracket and into position through the stays. It was then successfully located through into the cab with the help of Bill and Jon, and the regulator handle was fitted to avoid anyone bumping themselves on the protuding end of the rod. After lunch, the right hand connecting rod was fitted by the crosshead team: the little end has been tightened up and cotter pin put in but not opened. The big end brasses have been assembled on the rod and are fitted onto the journal but have not been fastened up as they may need shimming. It is sat on some timber for support.

The 12.40 departure for Pickering is hauled by Tornado -Nigel Hall

There were a lot of visitors, including our Treasurer, Richard Wheeler, but he would not accept the offer of a spare pair of overalls and to get stuck in - not even a cup of tea! - but did meet some of the MPD staff with whom he has regular dealings, but had not previously met. I had to leave about three o' clock as I had been invited to travel to Pickering and back on the A1 Trust special, and on which there were a number of NELPG members. Piglet came up to Deviation Shed about 4pm, just as the team was packing up, to work on Lucie for a few hours and he agreed to lock up the shed when he had finished.


Saturday 3rd March 2018


Ian reports another cold grey day with the odd snow flurry at Grosmont. All the main roads were open from Teesside and Wetherby with Ian McCall, Peter Ellis, Bill Dobson, Trevor Wilford and Ian in attendance. Around lunch time came Nigel Hall, followed by the hiab down Esk Valley bank with the J27 cab, chimney, sand boxes, springs and various other bits and pieces (see separate report and photo from last night).

The usual cuppa was prepared by Trevor (what a great guy) and no mice again. Work started with Trevor removing the last nut from the header, which was then removed, while Ian P, Ian M and Peter removed the left inside eccentric for drilling out the oil way hole which was too small for the syphon tube. After drilling and fitting the tube, this eccentric was fitted back onto the sheave with a 15 thou shim put in, and tightened up. Lunch was had in the warmth of the MPD mess room.

After lunch, the two Ian's had a quick look round Tornado. Then back to work trying to fit the left hand eccentric rod. This became very frustrating as it would not sit flush onto the eccentric. No matter how much we tried to knock it, and move it up and down, it did not want to move. It was taken down and some filing was done, put back up and still it would not fit. As time ran out, this is something for another go on Wednesday. Ian M started to make new cotters for the eccentrics, but didn't get far as he was helping Ian P and Bill with the eccentric rod.

Not much to report on the boiler: there's one or two chamfer's still to cut on the flues. Nigel delivered the Q6 valves. These are now stored behind the bench in Deviation Shed with a pink blanket over them.



Wednesday 28th February 2018


With no Q6 working party, there is nothing to report other than the damaged copper ferrule has been removed and Mark O'Brien is now sourcing some suitable copper bar to make a replacement. The remaining flue tubes await having the chamfers cut. Bill is chasing up Mathew Storey about progress with the welding of the elements.

  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  7 
  •  8 
  •  9 
  •  10 
  •  Next 
  •  End 
  • »

Page 1 of 28